### v1

@@ 4 Dataset
We introduce DataVisualWorkflow, a novel dataset for UI understanding in the context of computer workflows. The dataset is comprised of 14345 frames, where 13894 frames were extracted from 88 videos recorded by us and 451 were extracted from 10 videos downloaded from YouTube. DataVisualWorkflow includes 27 classes describing software and view and up to 81 classes describing the combination of software, view and context. It also contains labels for context menu and selected text recognition. The recorded videos are divided into train and test sets; the train set contains 9597 frames while the test set contains 4297 frames. The test set contains labels from software, view and context, whereas on the train set only software and view are labeled.
@@ 4.1 Dataset Recording
We acquired data from two different sources: mock-up workflows and YouTube. First, we recorded the mock-up workflows using OBS Studio’s [^1] https://obsproject.com/ screen recorder. The workflows were recorded by a single person over the course of a month at 25 FPS with a resolution of $1920\times 1080$. Most videos show long form coherent actions, like the ones presented on Figure 2. Others are composed of random short-form actions performed in a specific software application. Figure 3 shows the distribution of long-form action labels and videos with non-coherent long-form actions. As mentioned in Section 1 we have a specific interest in enterprise workflow automation, thus coherent long-form actions performed follow common enterprise use cases.
Later on, we downloaded video tutorials from YouTube following the constraint that long-form actions used to record the videos should be done in similar software applications to the ones in the recorded dataset.
4.2 Data Cleaning
Algorithm 1 Motion Detection Algorithm
function MotionDet($F[\;]$)
     $b$leftarrow $\varnothing$
     $a$leftarrow $\varnothing$
     for $i\leftarrow 1$ to $N$ do
         if $b=\varnothing$ then
              $b$leftarrow $F_{i}$
         else
              $a$leftarrow $F_{i}$
              $a_{diff}$leftarrow $\textsc{AbsDiff}(a,b,k_{G})$
              $c$leftarrow $\textsc{FindContours}(a_{diff},k_{D},T_{B})$
              $s$leftarrow $\textsc{CalcAreas}(a,b,c)$
              $b,a$leftarrow $\textsc{Save}(a,b,s,T_{C})$
         end if
     end for
end function
@@ 4.3 Data Labeling
At the time of recording, we used xdotool [^2] https://launchpad.net/ubuntu/trusty/+package/xdotool to record the name of the software application on screen and the view in the software. However, tools like xdotool are not perfect. They tend to make mistakes because they are out of sync with the recording software, names are not standardized, and some views just return an empty string. In order to have correct and standardized labels we had to do a thorough, frame by frame, quality assurance process. At the same time, we added the labels for context only on the test set and YouTube set. Table 1 shows the distribution of the software and view labels throughout the dataset, the distribution is consistent with enterprise use of computer software. Table 2 presents the distribution of the context labels in the test and YouTube sets.
Table 1: DataVisualWorkflow software-view labels distribution. Software Application View Percentage Recorded Percentage Youtube File Explorer Grid View 3.00 File Explorer Options 0.16 Web Browser Gmail 16.58 4.43 Web Browser Google 1.78 Web Browser Maps 6.58 Web Browser New Tab 2.27 Web Browser Save 0.22 Web Browser Web Page 2.74 Web Browser Web PDF 2.50 Web Browser Options 1.24 Spread Sheet Main View 24.85 42.35 Spread Sheet Save 0.21 Spread Sheet Options 0.30 5.23 Image Viewer Main View 0.69 Document Editor Main View 19.93 27.49 Document Editor Save 1.73 4.21 Document Editor Options 1.12 2.60 PDFViewer Main View 1.47 10.90 PDFViewer Save 0.9 PDFViewer Options 0.9 Terminal Main View 2.52 Mail Gmail 7.25 Mail Save 0.11 Desktop Main View 0.14 0.9 Presentation Editor Main View 2.32 Presentation Editor Save 0.02 Presentation Editor Options 0.30
@@
Label Percentage Recorded Percentage YouTube
Context Menu 7.26 10.20
Selected Text 30.93 39.02
None 61.81 50.77
[Uncaptioned image] Figure 3: Distribution of long-form actions and random actions in DataVisualWorkflow.
@@ 6.1 Datasets and Evaluation
We evaluate the model on our DataVisualWorkflow dataset. We train only on the recorded frames train set. The YouTube subset of our data is reserved to test generalization to out-of-distribution frames. This set contains different versions of software applications seen on the recorded set. It also contains applications not seen in the recorded set, but with the same basic functionality. Examples include Microsoft Word as an equivalent to Libre Office Writer, or Adobe PDF Reader as equivalent to Ubuntu’s PDFViewer.
@@
When testing recorded frames $svc$ classes, the test set is divided roughly in half by video id, so there is a low likelihood of 1-to-1 matching of frames in the database set and the query set. Due to the test set of the recorded frames being the only one with full labels, aside from the YouTube set, we use it as a database when testing OOD performance.
@@ 6.2 Model Configurations
Table 5: Generalization to OOD data results.
Method Loss Levels Emb. dim Software-View-Context
AMI R-Prec. mAP@R Prec@1
ST-Classification Cross Entropy 1 17.83
UISTCon SVC 1 $1\times 128$ 54.08 30.77 18.09 17.56
$1\times 256$ 53.54 25.91 16.71 23.74
$1\times 384$ 48.23 20.38 12.09 23.77
MT-Classification Cross Entropy 2 19.06
UISTCon SHL 2 $1\times 128$ 52.85 32.68 22.88 28.06
$1\times 256$ 50.86 22.26 14.23 19.85
$1\times 384$ 53.84 31.04 19.66 24.61
UISTCon HiConMulConE [58] 2 $1\times 128$ 50.53 25.31 16.60 14.95
$1\times 256$ 50.38 29.78 19.46 19.26
$1\times 384$ 50.45 30.60 20.42 26.55
UISTCon RINCE [17] 2 $1\times 128$ 44.85 26.76 16.19 19.99
$1\times 256$ 51.83 29.04 21.56 39.84
$1\times 384$ 50.19 23.50 14.73 14.81
UIMTCon SHL 2 $2\times 128$ 53.62 29.33 19.78 20.48
MT-Classification Cross Entropy 3 17.56
UISTCon SHL 3 $1\times 128$ 52.68 34.10 24.99 22.61
$1\times 256$ 49.05 27.42 17.34 22.57
$1\times 384$ 55.51 30.81 21.26 24.56
UISTCon HiConMulConE [58] 3 $1\times 128$ 52.97 21.00 13.34 17.94
$1\times 256$ 52.23 27.76 19.23 21.20
$1\times 384$ 57.05 28.43 18.78 20.58
UISTCon RINCE [17] 3 $1\times 128$ 43.51 15.60 8.47 16.84
$1\times 256$ 47.93 17.42 11.47 16.20
$1\times 384$ 49.08 21.07 14.88 15.56
UIMTCon SHL 3 $3\times 128$ 53.70 32.31 21.85 29.74
@@ 6.4 Results
Now, we analyze the performance of the models on the recorded frames and YouTube settings. Since there are no previous works solving this task, it is difficult to compare to previous works. Thus, our results show an ablation study and a comparison to previously proposed similar loss functions.
@@
We compare our proposed framework and baselines against supervised cross entropy on classification accuracy (precision@1). The same backbone, pre-trained weights, and training conditions were used. UIMTCon-2L scores the highest among all methods in precision@1, as shown in Table 3, while standard classification scored the lowest by a wide margin. Figure 7 further supports our results by showing strong classification performance of UIMTCon at high $k$ numbers for both UIMTCon-3L and UIMTCon-2L at $svc$ level, while UIMTCon-3L is able to perform comparatively to the best performing model at $s$ and $sv$ levels without $\text{FC}_{svc}$ being directly exposed to those labels. Additionally, by observing the results presented in Table 3, we can see that classification and retrieval metrics for RINCE [17], HiConMulConE [58] and SHL tend to go down or make small gains, compared to UIMTCon, when the software hierarchical level is added to the objective function. Experiments presented in Tables 7 and 6, in the supplementary material, show that this is likely due to noise introduced by the synthetic samples, this points to UIMTCon being robust to noise.
YouTube Frames: As stated before, computers can display massive changes in its UI from user to user. Therefore, it is important that models are able to keep their performance in this setting despite important appearance changes on irrelevant parts of the interface. All $s$ classes are found in the recorded set. However, some $sv$ labels are not present. Those are not taken into account to calculate the metrics found in Table 5.
Surprisingly, cross entropy models did not post results too inferior to models trained with contrastive loss when presented with OOD data. Experiments show that none of the losses tested are able to take into account such samples, even if they contain semantically similar classes. Table 5 shows that, in line with the recorded frames experiments, RINCE [17] and HiConMulConE [58] tend to lose performance when more hierarchical levels are used in the objective function, while the opposite happens with UIMTCon and SHL.
@@ Appendix 0.A Rest of Hierarchy Experiments
Table 6 and Table 8 show results at hierarchy levels $s$ and $sv$ on the recorded and YouTube respectively. As mentioned in the main paper, only $\text{FC}_{svc}$ is used to extract the representations used to calculate the metrics.
Experiments in Table 6 show UISTCon-3L has a competitive performance to the UISTCon-SHL at $sv$ level. Both UIMTCon-3L and UIMTCon-2L perform in line with most of other methods at the $s$ level at their respective hierarchy level usage group. Although UIMTCon posts unremarkable results in most metrics on the YouTube set (Table 8), we can observe that UIMTCon-3L scores third best in classification at the $sv$ hierarchy level.
@@ Appendix 0.C T-SNE Visualization
Visualizing the data is helpful to further understand the quantitative results. In order to visualize the results, we project the data into two dimensions using t-sne [33]. On the one hand, Figure 12 shows a more clear separation of classes in Figures 11(a) and 11(b), which coincides with the quantitative results shown in the main paper. On the other hand, Figure 13 illustrates that models trained without synthetic samples tend to generate clusters of a higher quality (Figures 12(d), 12(e) and 12(f)), when compared to the ones trained with synthetic samples (Figures 12(a), 12(b) and 12(c)).
@@ Appendix 0.F More Samples
Figure 9 shows samples taken at random from different videos of the recorded frames dataset. At the same time, in Figure 8 we show frames from the YouTube set of the DataVisualWorkflow dataset. Further sample sequences are illustrated in Figure 10 and Figure 11, which are sampled from the recorded frames and YouTube set respectively.
Figure 8: Sample images taken from DataVisualWorkflow’s YouTube set.
@@
Figure 11: Sample sequence taken from DataVisualWorkflow’s YouTube set.

### v2

@@ 4 Dataset
We introduce DataVisualWorkflow, a novel dataset for UI understanding in the context of computer workflows. The dataset is comprised of 13894 frames were extracted from 88 videos recorded by us. DataVisualWorkflow includes 26 classes describing software and view and up to 78 classes describing the combination of software, view and context. It also contains labels for context menu and selected text recognition. The recorded videos are divided into train and test sets; the train set contains 9597 frames while the test set contains 4297 frames. The test set contains labels from software, view and context, whereas on the train set only software and view are labeled.
@@ 4.1 Dataset Recording
We acquired data through mock-up workflows. First, we recorded the mock-up workflows using OBS Studio’s [^1] https://obsproject.com/ screen recorder. The workflows were recorded by a single person over the course of a month at 25 FPS with a resolution of $1920\times 1080$. Most videos show long form coherent actions, like the ones presented on Figure 2. Others are composed of random short-form actions performed in a specific software application. Figure 3 shows the distribution of long-form action labels and videos with non-coherent long-form actions. As mentioned in Section 1 we have a specific interest in enterprise workflow automation, thus coherent long-form actions performed follow common enterprise use cases.
@@ 4.2 Data Labeling
At the time of recording, we used xdotool [^2] https://launchpad.net/ubuntu/trusty/+package/xdotool to record the name of the software application on screen and the view in the software. However, tools like xdotool are not perfect. They tend to make mistakes because they are out of sync with the recording software, names are not standardized, and some views just return an empty string. In order to have correct and standardized labels we had to do a thorough, frame by frame, quality assurance process. At the same time, we added the labels for context only on the test set. Table 1 shows the distribution of the software and view labels throughout the dataset, the distribution is consistent with enterprise use of computer software. Table 2 presents the distribution of the context labels in the test.
Table 1: DataVisualWorkflow software-view labels distribution.
Software Application View Percentage Recorded
File Explorer Grid View 3.00
File Explorer Options 0.16
Web Browser Gmail 16.58
Web Browser Google 1.78
Web Browser Maps 6.58
Web Browser New Tab 2.27
Web Browser Save 0.22
Web Browser Web Page 2.74
Web Browser Web PDF 2.50
Web Browser Options 1.24
Spread Sheet Main View 24.85
Spread Sheet Save 0.21
Spread Sheet Options 0.30
Image Viewer Main View 0.69
Document Editor Main View 19.93
Document Editor Save 1.73
Document Editor Options 1.12
PDFViewer Main View 1.47
Terminal Main View 2.52
Mail Gmail 7.25
Mail Save 0.11
Desktop Main View 0.14
Presentation Editor Main View 2.32
Presentation Editor Save 0.02
Presentation Editor Options 0.30
@@
Label Percentage Recorded
Context Menu 7.26
Selected Text 30.93
None 61.81
Figure 3: Distribution of long-form actions and random actions in DataVisualWorkflow.
@@ 6.1 Datasets and Evaluation
We evaluate the model on our DataVisualWorkflow dataset. We train only on the pre-defined recorded frames train set.
@@
When testing recorded frames $svc$ classes, the test set is divided roughly in half by video id, so there is a low likelihood of 1-to-1 matching of frames in the database set and the query set.
@@ 6.4 Results
Now, we analyze the performance of the models on the DataVisualWorkflow. Since there are no previous works solving this task, it is difficult to compare to previous works. Thus, our results show an ablation study and a comparison to previously proposed similar loss functions.
@@
We compare our proposed framework and baselines against supervised cross entropy on classification accuracy (precision@1). The same backbone, pre-trained weights, and training conditions were used. UIMTCon-2L scores the highest among all methods in precision@1, as shown in Table 3, while standard classification scored the lowest by a wide margin. Figure 7 further supports our results by showing strong classification performance of UIMTCon at high $k$ numbers for both UIMTCon-3L and UIMTCon-2L at $svc$ level, while UIMTCon-3L is able to perform comparatively to the best performing model at $s$ and $sv$ levels without $\text{FC}_{svc}$ being directly exposed to those labels. Additionally, by observing the results presented in Table 3, we can see that classification and retrieval metrics for RINCE [17], HiConMulConE [58] and SHL tend to go down or make small gains, compared to UIMTCon, when the software hierarchical level is added to the objective function. Experiments presented in Tables 6 and 5, in the supplementary material, show that this is likely due to noise introduced by the synthetic samples, this points to UIMTCon being robust to noise.
@@ Appendix 0.A Rest of Hierarchy Experiments
Table 5 and Table 7 show results at hierarchy levels $s$ and $sv$. As mentioned in the main paper, only $\text{FC}_{svc}$ is used to extract the representations used to calculate the metrics.
Experiments in Table 5 show UISTCon-3L has a competitive performance to the UISTCon-SHL at $sv$ level. Both UIMTCon-3L and UIMTCon-2L perform in line with most of other methods at the $s$ level at their respective hierarchy level usage group.
@@ Appendix 0.C T-SNE Visualization
Visualizing the data is helpful to further understand the quantitative results. In order to visualize the results, we project the data into two dimensions using t-sne [33]. On the one hand, Figure 10 shows a more clear separation of classes in Figures 10(a) and 10(b), which coincides with the quantitative results shown in the main paper. On the other hand, Figure 11 illustrates that models trained without synthetic samples tend to generate clusters of a higher quality (Figures 11(d), 11(e) and 11(f)), when compared to the ones trained with synthetic samples (Figures 11(a), 11(b) and 11(c)).
@@ Appendix 0.F More Samples
Figure 8 shows samples taken at random from different videos of the recorded frames dataset. Sample sequences are illustrated in Figure 9.
@@
Appendix 0.G Data Cleaning
Algorithm 1 Motion Detection Algorithm
function MotionDet($F[\;]$)
     $b$leftarrow $\varnothing$
     $a$leftarrow $\varnothing$
     for $i\leftarrow 1$ to $N$ do
         if $b=\varnothing$ then
              $b$leftarrow $F_{i}$
         else
              $a$leftarrow $F_{i}$
              $a_{diff}$leftarrow $\textsc{AbsDiff}(a,b,k_{G})$
              $c$leftarrow $\textsc{FindContours}(a_{diff},k_{D},T_{B})$
              $s$leftarrow $\textsc{CalcAreas}(a,b,c)$
              $b,a$leftarrow $\textsc{Save}(a,b,s,T_{C})$
         end if
     end for
end function